# Greedy Perspectives: Multi-Drone View Planning for Collaborative Perception in Cluttered Environments

### v2

@@ Abstract
Deployment of teams of aerial robots could enable large-scale filming of dynamic groups of people (actors) in complex environments for applications in areas such as team sports and cinematography. Toward this end, methods for submodular maximization via sequential greedy planning can be used for scalable optimization of camera views across teams of robots but face challenges with efficient coordination in cluttered environments. Obstacles can produce occlusions and increase chances of inter-robot collision which can violate requirements for near-optimality guarantees. To coordinate teams of aerial robots in filming groups of people in dense environments, a more general view-planning approach is required. We explore how collision and occlusion impact performance in filming applications through the development of a multi-robot multi-actor view planner with an occlusion-aware objective for filming groups of people and compare with a formation planner and a greedy planner that ignores inter-robot collisions. We evaluate our approach based on five test environments and complex multi-actor behaviors. Compared with a formation planner, our sequential planner generates 14% greater view reward over the actors for three scenarios and comparable performance to formation planning on two others. We also observe near identical view rewards for sequential planning both with and without inter-robot collision constraints which indicates that robots are able to avoid collisions without impairing performance in the perception task. Overall, we demonstrate effective coordination of teams of aerial robots for filming groups that may split, merge, or spread apart and in environments cluttered with obstacles that may cause collisions or occlusions.
@@ I Introduction
The capture of significant events via photos and video has become universal, and Unmanned aerial vehicles (UAVs) extend the capabilities of cameras by allowing for view placement in otherwise hard-to-reach places and tracking intricate trajectories. Multiple aerial cameras can be used to not only view an actor from multiple angles simultaneously but perform higher functions such as localization and tracking [1, 2, 3], environment exploration and mapping [4, 3], cinematic filming [5, 6, 7, 8], and outdoor human pose reconstruction [9, 10]. These applications rely on effective collaboration between groups of UAVs whereas manual control may result in poor shot selection and view duplication while requiring many coordinated operators. Therefore, autonomous coordination of UAV teams may be necessary for tasks such as multi-robot filming or reconstruction. However, directly maximizing domain-specific metrics, such as reconstruction accuracy, can be difficult to perform online—this motivates development of proxy objectives that quantify coverage and detail for multiple views. For example, Bucker et al. [5] demonstrate cinematic filming through a joint objective combining collision and occlusion avoidance, shot diversity, and artistic principles in filming a single actor. We are interested in similar settings but where robots collaborate to obtain diverse views of multiple actors—in a cluttered environment, with occlusions, where robots may observe more than one actor at once.
@@
While defining an objective can be difficult, planning for multi-robot aerial systems also presents a challenge; the vast joint state space, non-convex environment, and non-linear view rewards make optimal planning intractable. Many applications exploit problem-specific structures to reduce the overall such as by optimizing an actor-centric formation [9, 12] or by altering alter the search procedure to generate single-robot trajectories sequentially [5, 4, 1, 2, 3]. In this work, we apply a planning approach much like Bucker et al. [5], and develop a system design and view rewards that enable application to a multi-actor setting.
Figure 2: View Planning System Overview The multi-actor scenario is translated to an internal planner representation. The Markov Decision Process with DAG structure encodes collision constraints and view rewards. The Multi-Robot View Plan is produced through sequential greedy planning. Joint multi-drone trajectories are brought back to continuous 3D coordinates and output to the navigation stack.
Problem: The dynamic multi-actor view planning problem consists of generating sequences of camera views over a fixed planning horizon to maximize a collective view reward which is a function of pixel densities over the surfaces of the actors. The primary assumptions for our approach are: known static environment, known actor trajectories (e.g. scripted scenarios), and known robot start state. An illustration of the problem setup is depicted in Fig. 1. These assumptions are fairly strict for the purpose of evaluating the planning approach. In practice, our approach could be applied in a receding-horizon setting based on scripted or predicted actor trajectories [5].
@@
  • An ccclusion-aware objective based on rendered camera views for filming groups of people
@@ II Related Works
Aerial Filming: Aerial perception systems have grown to widespread use through their success in providing low-cost filming of conventionally challenging unscripted scenes. Consumer and commercial systems such as the Skydio S2+ [15] demonstrate single-drone filming capabilities and are starting to incorporate collaborative multi-drone behaviors for mapping. Research developing autonomous aerial filming systems for cinematography has also focused on developing systems that can perform parametrized actions such as tracking a subject, rule of thirds framing, or dolly zoom by methods such as model predictive control to optimize motion and camera controls [7, 8] and by scheduling multi-robot missions [16, 17, 18, 19]. Another body of work focuses on learning artistic principles rather than implementing them by hand [20, 21]. The focus of these works on mission planning and implementing and learning artistic principles is complementary to our approach which develops a capability for filming complex group behaviors with occlusions.
@@
Sequential and Submodular Multi-Robot Planning: Typically multi-robot perception planning and information gathering problems, cannot tractably be solved optimally due to their combinatorial nature, but greedy methods for submodular optimization can often promise information gain or perception quality no worse than half of optimal in polynomial time [26, 27]. Submodular optimization and sequential greedy planning has been applied extensively to such multi-robot coordination problems [28, 4, 1, 2, 3, 29, 30, 5]. However, questions of occlusions and camera views have been explored primarily in the setting of mapping and exploration [4, 3]. Lauri et al. [29] present an exception in which eye-in-hand cameras inspect and reconstruct static scenes. Unlike exploration and mapping, applications involving filming moving actors can force persistent interaction between robots over the duration of a planning horizon or experiment, and our early work on this topic indicates that sequential planning is important for effective cooperation in settings involving observing moving subjects [31].
@@ III-A Submodularity and Monotonicity
The view reward we employ satisfies monotonicity properties that are useful in developing our approach to planning and coordination. Informally, submodularity expresses the principle of diminishing returns and monotonicity requires functions to be always increasing. Given a set of actions $\Omega$, a set function $g$ maps subsets of actions (robots’ plans) to a real value (the view reward). A set function is monotonic if adding elements to a set does not decrease its value; that is for any $A\subseteq B\subseteq\Omega$ then $g(A)\leq g(B)$. A set function is submodular if marginal gains decrease monotonically; specifically, given $A\subseteq B\subseteq\Omega$ and $C\subseteq\Omega\setminus B$, then $g(A\cup C)-g(A)\geq g(B\cup C)-g(B)$. Objectives related to perception planning [32, 29] and information gathering [28] are often submodular, and this corresponds to how marginal view rewards may diminish given by repeated views of the same actor from the same perspective.
@@ III-B Partition Matroids
A partition matroid can be used to represent product-spaces of actions or trajectories that arise in multi-robot planning problems [33, Sec. 39.4]. Consider a view planning problem involving a team of robots =$\mathcal{R}=\{1,\ldots,N^{\mathrm{r}}\}$ where each robot $i\in\mathcal{R}$ has access to a set of actions $\mathscr{U}_{i}$. These actions can take many forms such as assignments, trajectories, or paths. The set of all actions for a robot is the ground set $\Omega=\bigcup_{i\in\mathcal{R}}\mathscr{U}_{i}$. Each robot is assigned one action from its corresponding set $\mathscr{U}_{i}$. If there are no collisions between robots the set of valid and partial assignments forms a partition matroid: =$\mathscr{I}=\{X\subseteq\Omega\mid 1\geq|X\cap\mathscr{U}_{i}|\ \forall i\in\mathcal{R}\}$. To satisfy this structure each robots’ actions must be interchangeable to satisfy the exchange property of a matroid. Inter-robot collisions violate this property because swapping actions can cause conflicts with other robots.
@@ IV Problem Formulation
We aim to coordinate a team of UAVs to maximize coverage-like view rewards for observing a group of actors moving through an obstacle-dense environment. Consider a set of actors $\mathcal{A}=\{1,\ldots,N^{\mathrm{a}}\}$ each with a set of faces $\mathcal{F}_{a}=\{1,\ldots,N^{\mathrm{f}}_{a}\}$ where $a\in\mathcal{A}$ and a set of robots =$\mathcal{R}=\{1,\ldots,N^{\mathrm{r}}\}$. Each robot $i\in\mathcal{R}$ can execute a control action $u_{i,t}\in\mathcal{U}_{i}\subseteq SE(2)$ at time $t\in\{0,\ldots,T\}$. The robots go on to select a finite-horizon sequence of viable control actions which form their plans. Additionally, robots have associated states $x_{i,t}\in\mathcal{X}$ which is a subset of $SE(2)$. Sequences of states form the robots’ trajectories $\xi_{i}=[x_{i,0},\ldots,x_{i,T}]$—we will occasionally index trajectories to obtain $\xi_{i,t}=x_{i,t}$. Each trajectory, once fixed, produces non-collision constraints for all other robots. Given the trajectories of all actors in $SE(3)$, start states $x_{i,0}$, and environment geometry, we aim to find joint collision-free control sequences $U^{\ast}=\bigcup_{i\in\mathcal{R}}[u_{i,0},\ldots,u_{i,T}]$ that maximize our objective and fit our motion model.
@@ IV-B Non-Collision Constraints
Robots are considered in collision with the environment when the discretized state location is occupied by an element of the environment map that exceeds the robot’s height. Similarly, a pair of robots is in collision when both occupy the same discretized cell at the same time. We implicitly assume a conservative model of the environment (the heightmap and discretization of the state space) to ensure safety of states that satisfy the obstacle and inter-robot collision constraints.
@@ IV-C Camera and View Reward Model
In order to encourage robots to assume views that uniformly cover the actors and their corresponding faces, we apply a square root to introduce diminishing returns on increasing pixel densities from multiple views [14, 34]. Finally, given robot trajectories according to the dynamics (1) and selected control inputs, the robots obtain the following view reward for the given face and time:
@@ IV-D Objective
So, robot $i\in\mathcal{R}$ obtains $\sum_{t\in\{0,\ldots,T\}}R_{s}(u_{i,t})$ reward for timesteps it remains stationary. The joint objective is then as follows:
@@ V-A Evaluation of View Reward
To produce an occlusion-aware implementation of our view reward (2) we compute pixels by implementing an OpenGL rasterization renderer based on a 2.5D height map of the environment and simplified actor geometry—we use polygonal cylinders. We then use a perspective camera based on specified camera intrinsics to capture an occlusion-aware representation of the scene from a given robot state. The system renders the environment via the GPU with a geometry shader to draw the heightmap. To determine how many pixels the cameras observe for each face, we render the faces with unique colors associated with the actor and face IDs. Counting pixels of each color and dividing by the areas of the faces yields the corresponding pixel densities. Figure 3 illustrates this process and provides an example of a rendered view.
@@ V-C Sequential Planning
Now, we are able to generate the joint view plans for the multi-robot team. We do so by sequentially planning greedy single-robot trajectories as is common for methods based on submodular optimization [28, 2, 5, 3, 1, 4, 29, 30]. With some abuse of notation, each robot maximizes the objective for itself with access to prior decisions in the sequence:
@@
This forms a series of single-robot subproblems that we solve with the value iteration planner (Sec. V-B). Through the course of this process, we accumulate pixel densities per each face to evaluate the view reward (2) and filter out states that would produce collisions with other robots (Sec. IV-B).
@@ VI-A Formation Planning
We compare our sequential view planner against an assignment and formation based planner which we model off the multi-view formations applied in [12, 9] following analysis by Bishop et al. [37]. We assign equal numbers of robots to each actor, and groups assume formations as follows. The formation has a constant radius around an actor and a separation angle $\phi$; for $N^{\mathrm{r}}>2$ then $\phi=\frac{2\pi}{N^{\mathrm{r}}}$ and when $N^{\mathrm{r}}=2$ then $\phi=\frac{\pi}{2}$ (see Fig. 5). We directly orient the formation about the actor to maximize the view reward $R_{\mathrm{v}}$ (including all actors) at each timestep. The formation planner ignores the motion model (Sec. IV-A) and does not consider environment and robot collisions; so view rewards for formation planning are generally optimistic.
@@ VI-C Sequential Planner Performance
Figure 7: Average view reward normalized by the baseline formation planner performance. 10 unique robot start configurations were specified for the sequential planners and are compared with the unique output from the formation planner. Both sequential planners outperform the formation planner baseline in the Merge, Corridor, and Forest scenarios, or else perform similarly. Both versions of the sequential planner perform similarly in all scenarios; view reward for the collision-aware planner is not impaired, though that planner no longer satisfies guarantees on solution quality.
@@ VI-D Scaling number of robots
Figure 8: Compute time and reward accumulated with scaling the number of robots in the Large test case. The means and standard deviations are computed across 10 unique start locations.

### v3

@@ Abstract
Deployment of teams of aerial robots could enable large-scale filming of dynamic groups of people (actors) in complex environments for applications in areas such as team sports and cinematography. Toward this end, methods for submodular maximization via sequential greedy planning can enable scalable optimization of camera views across teams of robots but face challenges with efficient coordination in cluttered environments. Obstacles can produce occlusions and increase chances of inter-robot collision which can violate requirements for near-optimality guarantees. To coordinate teams of aerial robots in filming groups of people in dense environments, a more general view-planning approach is required. We explore how collision and occlusion impact performance in filming applications through the development of a multi-robot multi-actor view planner with an occlusion-aware objective for filming groups of people and compare with a formation planner and a greedy planner that ignores inter-robot collisions. We evaluate our approach based on five test environments and complex multi-actor behaviors. Compared with a formation planner, our sequential planner generates 14% greater view reward for filming the actors in three scenarios and comparable performance to formation planning on two others. We also observe near identical view rewards for sequential planning both with and without inter-robot collision constraints which indicates that robots are able to avoid collisions without impairing performance in the perception task. Overall, we demonstrate effective coordination of teams of aerial robots in environments cluttered with obstacles that may cause collisions or occlusions and for filming groups that may split, merge, or spread apart. Our implementation and the data used to produce results for this paper are available via the companion website: https://greedyperspectives.github.io/
@@ I Introduction
The capture of significant events via photos and video has become universal, and Unmanned aerial vehicles (UAVs) extend the capabilities of cameras by allowing for view placement in otherwise hard-to-reach places and tracking intricate trajectories. Multiple aerial cameras can be used to not only view an actor from multiple angles simultaneously but perform higher functions such as localization and tracking [1, 2, 3, 4, 5], environment exploration and mapping [6, 3], cinematic filming [7, 8, 9, 10], and outdoor human pose reconstruction [11, 12]. These applications rely on effective collaboration between groups of UAVs whereas manual control may result in poor shot selection and view duplication while requiring many coordinated operators. Therefore, autonomous coordination of UAV teams may be necessary for tasks such as multi-robot filming or reconstruction. However, directly maximizing domain-specific metrics, such as reconstruction accuracy, can be difficult to perform online—this motivates development of proxy objectives that quantify coverage and detail for multiple views. For example, Bucker et al. [7] demonstrate cinematic filming through a joint objective combining collision and occlusion avoidance, shot diversity, and artistic principles in filming a single actor. We are interested in similar settings but where robots collaborate to obtain diverse views of multiple actors—in a cluttered environment, with occlusions, where robots may observe multiple actors at once.
@@
While defining an objective can be difficult, planning for multi-robot aerial systems also presents a challenge: the vast joint state space, non-convex environment, and non-linear view rewards make optimal planning intractable. Many applications exploit problem-specific structures to reduce the problem complexity such as by optimizing an actor-centric formation [11, 14, 4] or by altering the search procedure to generate single-robot trajectories sequentially [7, 6, 1, 2, 3]. In this work, we apply a planning approach much like Bucker et al. [7], and develop a system design and view rewards that enable application to a multi-actor setting.
Figure 2: View Planning System Overview The multi-actor scenario is translated to an internal planner representation. The Markov Decision Process with DAG structure encodes collision constraints and view rewards. The Multi-Robot View Plan is produced through sequential greedy planning.
Problem: The dynamic multi-actor view planning problem consists of generating sequences of camera views over a fixed planning horizon to maximize a collective view reward that is a function of pixel densities over the surfaces of the actors. The primary assumptions for our approach are: known static environment, known actor trajectories (e.g. scripted scenarios), and known robot start state. An illustration of the problem setup is depicted in Fig. 1. These assumptions are fairly strict for the purpose of evaluating the planning approach. In practice, our approach could be applied in a receding-horizon setting based on scripted or predicted actor trajectories [7].
@@
  • An occlusion-aware objective based on rendered camera views for filming groups of people
@@ II Related Works
Aerial Filming: Aerial perception systems have grown to widespread use through their success in providing low-cost filming of conventionally challenging unscripted scenes. Consumer and commercial systems such as the Skydio S2+ [17] demonstrate single-drone filming capabilities and are starting to incorporate collaborative multi-drone behaviors for mapping. Research developing autonomous aerial filming systems for cinematography has also focused on developing systems that can perform parametrized actions such as tracking a subject, rule of thirds framing, or dolly zoom by methods such as model predictive control to optimize motion and camera controls [9, 10] and by scheduling multi-robot missions [18, 19, 20, 21]. Another body of work focuses on learning artistic principles rather than implementing them by hand [22, 23]. The focus of these works on mission planning and learning artistic principles is complementary to our approach which develops a capability for filming complex group behaviors with occlusions.
@@
Target tracking and localization: The filming scenario we study is similar in composition to target localization and tracking problems that focus on estimating position or motion of one or more targets [1, 2, 3, 4, 5]. While these works sometimes reason about field-of-view and occlusions [4] this reasoning is often limited and secondary to the tracking task—a task that is often formulated in terms of minimizing uncertainty in target state via an information-theoretic objective [1, 2, 3, 5]. Target tracking may also be thought of as representing a subsystem of a hypothetical filming system that relaxes our requirement for known trajectories—from this perspective, we observe that several works in this area apply sequential and auction-based methods for coordination that are generally compatible with our approach [1, 2, 3, 5].
Sequential and Submodular Multi-Robot Planning: Typically multi-robot perception planning and information gathering problems, cannot tractably be solved optimally due to their combinatorial nature, but greedy methods for submodular optimization can often promise information gain or perception quality no worse than half of optimal in polynomial time [28, 29]. Submodular optimization and sequential greedy planning has been applied extensively to such multi-robot coordination problems [30, 6, 1, 2, 3, 31, 32, 7]. However, questions of occlusions and camera views have been explored primarily in the setting of mapping and exploration [6, 3]. Lauri et al. [31] present an exception in which eye-in-hand cameras inspect and reconstruct static scenes. Unlike exploration and mapping, applications involving filming moving actors can force persistent interaction between robots over the duration of a scenario or planning horizon, and our early work on this topic indicates that sequential planning is important for effective cooperation in settings involving observing moving subjects [33].
@@ III-A Submodularity and Monotonicity
The view reward we employ satisfies monotonicity properties that are useful in developing our approach to planning and coordination. Informally, submodularity expresses the principle of diminishing returns and monotonicity requiring functions to be always increasing. Given a set of actions $\Omega$, a set function $g$ maps subsets of actions (robots’ plans) to a real value (the view reward). A set function is monotonic if adding elements to a set does not decrease its value; that is for any $A\subseteq B\subseteq\Omega$ then $g(A)\leq g(B)$. A set function is submodular if marginal gains decrease monotonically; specifically, given $A\subseteq B\subseteq\Omega$ and $C\subseteq\Omega\setminus B$, then $g(A\cup C)-g(A)\geq g(B\cup C)-g(B)$. Objectives related to perception planning [34, 31] and information gathering [30] are often submodular, and this corresponds to how marginal view rewards may diminish given by repeated views of the same actor from the same perspective.
@@ III-B Partition Matroids
A partition matroid can be used to represent product-spaces of actions or trajectories that arise in multi-robot planning problems [35, Sec. 39.4]. Consider a view planning problem involving a team of robots =$\mathcal{R}=\{1,\ldots,N^{\mathrm{r}}\}$ where each robot $i\in\mathcal{R}$ has access to a set of actions $\mathscr{U}_{i}$. These actions can take many forms such as assignments, trajectories, or paths. The set of all actions for a robot is the ground set $\Omega=\bigcup_{i\in\mathcal{R}}\mathscr{U}_{i}$. Each robot is assigned one action from its corresponding set $\mathscr{U}_{i}$. If there are no collisions between robots the set of valid and partial assignments forms a partition matroid: =$\mathscr{I}=\{X\subseteq\Omega\mid 1\geq|X\cap\mathscr{U}_{i}|\ \forall i\in\mathcal{R}\}$. To satisfy this structure each robots’ actions must be interchangeable to satisfy the exchange property of a matroid. Inter-robot collisions violate this property because swapping actions can cause conflicts with other robots.
@@ IV Problem Formulation
We aim to coordinate a team of UAVs to maximize coverage-like view rewards for observing a group of actors moving through an obstacle-dense environment. Consider a set of actors $\mathcal{A}=\{1,\ldots,N^{\mathrm{a}}\}$ each with a set of faces $\mathcal{F}_{a}=\{1,\ldots,N^{\mathrm{f}}_{a}\}$ where $a\in\mathcal{A}$ and a set of robots =$\mathcal{R}=\{1,\ldots,N^{\mathrm{r}}\}$. Each robot $i\in\mathcal{R}$ can execute a control action $u_{i,t}\in\mathcal{U}_{i}\subseteq SE(2)$ at time $t\in\{0,\ldots,T\}$. The robots go on to select a finite-horizon sequence of viable control actions that form their plans. Additionally, robots have associated states $x_{i,t}\in\mathcal{X}$ which is a subset of $SE(2)$. Sequences of states form the robots’ trajectories $\xi_{i}=[x_{i,0},\ldots,x_{i,T}]$—we will occasionally index trajectories to obtain $\xi_{i,t}=x_{i,t}$. Each trajectory, once fixed, produces non-collision constraints for all other robots. Given the trajectories of all actors in $SE(3)$, start states $x_{i,0}$, and environment geometry, we aim to find joint collision-free control sequences $U^{\ast}=\bigcup_{i\in\mathcal{R}}[u_{i,0},\ldots,u_{i,T}]$ that maximize our objective and fit our motion model.
@@ IV-B Non-Collision Constraints
Robots are considered in collision with the environment when the discretized state location is occupied by an element of the environment map that exceeds the robot’s height. Similarly, a pair of robots is in collision when both occupy the same discretized cell at the same time. We implicitly assume a conservative model of the environment (the height-map and discretization of the state space) to ensure safety of states that satisfy the obstacle and inter-robot collision constraints.
@@ IV-C Camera and View Reward Model
In order to encourage robots to assume views that uniformly cover the actors and their corresponding faces, we apply a square root to introduce diminishing returns on increasing pixel densities from multiple views [16]. Finally, given robot trajectories according to the dynamics (1) and selected control inputs, the robots obtain the following view reward for the given face and time:
@@ IV-D Objective
So, robot $i\in\mathcal{R}$ obtains $\sum_{t\in\{0,\ldots,T\}}R_{s}(u_{i,t})$ reward for time-steps it remains stationary. The joint objective is then as follows:
@@ V-A Evaluation of View Reward
To produce an occlusion-aware implementation of our view reward (2) we compute pixels by implementing an OpenGL rasterization renderer based on a 2.5D height map of the environment and simplified actor geometry—we use polygonal cylinders. We then use a perspective camera based on specified camera intrinsics to capture an occlusion-aware representation of the scene from a given robot state. The system renders the environment via the GPU with a geometry shader to draw the height-map. To determine how many pixels the cameras observe for each face, we render the faces with unique colors associated with the actor and face IDs. Counting pixels of each color and dividing by the areas of the faces yields the corresponding pixel densities. Figure 3 illustrates this process and provides an example of a rendered view.
@@ V-C Sequential Planning
Now, we are able to generate the joint view plans for the multi-robot team. We do so by sequentially planning greedy single-robot trajectories in an arbitrary order as is common for methods based on submodular optimization [30, 2, 7, 3, 1, 6, 31, 32]. With some abuse of notation, each robot maximizes the objective for itself with access to prior decisions in the sequence:
@@
This forms a series of single-robot sub-problems that we solve with the value iteration planner (Sec. V-B). Through the course of this process, we accumulate pixel densities per each face to evaluate the view reward (2) and filter out states that would produce collisions with other robots (Sec. IV-B).
@@ VI-A Formation Planning
We compare our sequential view planner against an assignment and formation based planner which we model off the multi-view formations applied in [14, 11] following analysis by Bishop et al. [38]. We assign equal numbers of robots to each actor, and groups assume formations as follows. The formation has a constant radius around an actor and a separation angle $\phi$; for $N^{\mathrm{r}}>2$ then $\phi=\frac{2\pi}{N^{\mathrm{r}}}$ and when $N^{\mathrm{r}}=2$ then $\phi=\frac{\pi}{2}$ (see Fig. 5). We directly orient the formation about the actor to maximize the view reward $R_{\mathrm{v}}$ (including all actors) at each time-step. The formation planner has a fixed radius, ignores the motion model (Sec. IV-A), and does not consider environment and robot collisions; so view rewards for formation planning are generally optimistic and require no-extra computation.
@@ VI-C Sequential Planner Performance
Figure 7: Average view reward normalized by the baseline formation planner performance. 10 unique robot start configurations were specified for the sequential planners and are compared with the unique output from the formation planner. Both sequential planners outperform the formation planner baseline in the Merge, Corridor, and Forest scenarios, or else perform similarly. Sequential planners (with and without collision avoidance) also behave similarly in terms of view rewards: view rewards for the collision-aware planner are not impaired, though that planner no longer satisfies guarantees on solution quality.
@@ VI-D Scaling number of robots
Figure 8: Computation time and reward accumulated with scaling the number of robots in the Large test case. The means and standard deviations are computed across 10 unique start locations.